# Emergent spin and orbital angular momentum of light in twisted photonic bilayer

### v1

@@ III.2 Moiré diffracted beams
The emergence of moiré pattern in twisted bilayers gives rise to a series of propagating diffracted beams with the in-plane wave vectors $|\bm{g}|<q_{0}$. To the lowest order in the permittivity modulation, there are four forward diffracted beams with the in-plane wave vectors ±$\pm\bm{K}$ and ±$\pm\bm{K}^{\prime}$, where $\bm{K}=\bm{k}_{1}-\bm{k}_{2}$ and $\bm{K}^{\prime}=\bm{k}_{1}^{\prime}-\bm{k}_{2}^{\prime}$. In the bilayers with small twist angles $\varphi\ll\omega/(ck)$, these beams slightly deviate from the incident beam direction.
@@
Figure 3(b) shows the intensity of the diffracted beams as a function of the interlayer distance and light frequency. The moiré diffraction stems from the near field interaction and, therefore, is highly sensitive to the interlayer distance. The diffraction is most prominent at the waveguide resonances. The SAM and OAM are shown in Figs. 3(c) and 3(d). The diffracted beam acquires much larger spin and orbital angular momenta than the transmitted or reflected beams. This is especially noticeable for the SAM, Fig. 3(c).
In the case of small layer polarizability, $\alpha_{0}k,\alpha_{1}k\ll 1$, the analytic expression for amplitudes of the moiré diffracted beams can be obtained, see Appendix B. For small twist angles, $\varphi\ll 1$, the SAM and OAM assume the form
@@
The SAM and OAM in the diffracted beams are much larger than those in the transmitted and reflected beams since the very occurrence of diffracted beams originates from the structure twist.
III.3 Field distribution in interference zone
Figure 4: (a) The spatial distribution of the SAM behind the bilayer in the area of interference of transmitted and diffracted beams. The incident beam is linearly polarized, $\beta=\pi/4$, $\varphi=\pi/15$, $kd=2.5$, $\omega/(ck)=0.7$. Other parameters are listed in the caption to Fig. 2. (b) The SAM at the coordinate $\bm{\rho}=[\pi/(k\varphi),0]$, highlighted by the green point in Fig. 4(a), as a function of the interlayer distance and the light frequency.
@@ IV Summary
The interaction of light with twisted photonic bilayer includes the transfer of angular momentum. As a result, initially unpolarized beam carrying no orbital angular momentum acquires spin angular momentum upon transmission and orbital angular momentum upon reflection. The effects are particularly pronounced at the frequencies of the eigen optical modes of the twisted bilayer. Both effects are sensitive to the interlayer distance but exhibit different behavior: the orbital angular momentum in reflection oscillates with a constant amplitude when the interlayer distance is increased while the spin angular momentum in transmission decays exponentially at large distances. The spin angular momentum emerges from the near field interaction in the twisted bilayer and the related diffraction by the moiré pattern. The origin of the orbital angular momentum is the interference of the asymmetric distortions of the beam induced by the two layers. The moiré diffracted beams are also found to carry significant spin and orbital angular momenta.

### v2

@@ III.3 Moiré diffracted beams
The emergence of moiré pattern in twisted bilayers gives rise to a series of propagating diffracted beams with the in-plane wave vectors $|\bm{g}|<q_{0}$. To the lowest order in the permittivity modulation, there are four forward and four backward diffracted beams with the in-plane wave vectors ±$\pm\bm{K}$ and ±$\pm\bm{K}^{\prime}$, where $\bm{K}=\bm{k}_{1}-\bm{k}_{2}$ and $\bm{K}^{\prime}=\bm{k}_{1}^{\prime}-\bm{k}_{2}^{\prime}$. In the bilayers with small twist angles $\varphi\ll\omega/(ck)$, these beams slightly deviate from the bilayer normal.
@@
Figure 5(b) shows the intensity of the forward diffracted beam as a function of the interlayer distance and light frequency. The moiré diffraction stems from the near-field interaction and, therefore, is highly sensitive to the interlayer distance and the twist angle. The diffraction is most prominent at the waveguide resonances. The SAM and OAM of the diffracted beam are shown in Figs. 5(c) and 5(d). The diffracted beam acquires much larger spin and orbital angular momenta than the transmitted or reflected beams. This is especially noticeable for the SAM, Fig. 5(c).
In the case of small layer polarizability, $\alpha_{0}k,\alpha_{1}k\ll 1$, the analytic expression for amplitudes of the moiré diffracted beams can be obtained, see Appendix B. For small twist angles, $\varphi\ll 1$, the intensity of each forward diffracted beam is given by
$T^{(\bm{K})}=\frac{q_{0}^{2}}{\varkappa^{2}}(2\pi\alpha_{1})^{4}(\varkappa^{4}+q_{0}^{4})[1+\cos(2q_{0}d)]{\rm e}^{-2\varkappa d}\;,$ (8)
while the SAM and OAM assume the form
@@
Both $\langle s\rangle$ and $\langle l\rangle$ in Eqs. (9) and (10), respectively, are independent of the permittivity modulation $\alpha_{1}$. The reason is that the very occurrence of the moiré diffracted beams originates from the structure twist. The fluxes of the SAM and OAM as well as the photon flux are all proportional to $\alpha_{1}^{4}$, see Appendix B. As a result, the SAM and OAM per photon are independent of the permittivity modulation and can reach high values.
III.4 Field distribution in interference zone
Figure 6: (a) The spatial distribution of the SAM behind the bilayer in the area of interference of transmitted and diffracted beams. The incident beam is linearly polarized, $\beta=\pi/4$, $\varphi=\pi/15$, $kd=2$, $\omega/(ck)=0.816$. Other parameters are listed in the caption to Fig. 2. (b) The SAM at the coordinate $\bm{\rho}=[\pi/(k\varphi),0]$, highlighted by the green point in Fig. 6(a), as a function of the interlayer distance and the light frequency.
@@ IV Summary
The interaction of light with twisted photonic bilayer includes the transfer of angular momentum. As a result, initially unpolarized beam carrying no orbital angular momentum acquires spin angular momentum upon transmission and orbital angular momentum upon reflection. The effects are particularly pronounced at the frequencies of the eigen optical modes of the twisted bilayer. Both effects are sensitive to the interlayer distance but exhibit different behavior: the orbital angular momentum in reflection oscillates with a constant amplitude when the interlayer distance is increased while the spin angular momentum in transmission oscillates and decays exponentially at large distances. The spin angular momentum emerges from the near-field interaction in the twisted bilayer and the related diffraction by the moiré pattern. The origin of the orbital angular momentum is the interference of the asymmetric distortions of the beam induced by the two layers. The moiré diffracted beams are also found to carry significant spin and orbital angular momenta. The predicted effects can be observed in twisted photonic-crystal bilayers which are already available in the infrared [38] and microwave [40] spectral ranges.